Getting Started with RTEMS

Edition 4.9.1, for 4.9.1 12 December 2008

On-Line Applications Research Corporation

On-Line Applications Research Corporation TEXinfo 2007-12-02.17

COPYRIGHT \begin{matrix}c \\ ◯ \end{matrix} 1988 - 2008. On-Line Applications Research Corporation (OAR).

The authors have used their best efforts in preparing this material. These efforts include the development, research, and testing of the theories and programs to determine their effectiveness. No warranty of any kind, expressed or implied, with regard to the software or the material contained in this document is provided. No liability arising out of the application or use of any product described in this document is assumed. The authors reserve the right to revise this material and to make changes from time to time in the content hereof without obligation to notify anyone of such revision or changes.

The RTEMS Project is hosted at http://www.rtems.com. Any inquiries concerning RTEMS, its related support components, its documentation, or any custom services for RTEMS should be directed to the contacts listed on that site. A current list of RTEMS

Support Providers is at http://www.rtems.com/support.html.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Real-Time Embedded Systems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Cross Development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Resources on the Internet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3.1 Online Tool Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.2 RTEMS Mailing List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.3 CrossGCC Mailing List. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.4 GCC Mailing Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 Disk Space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 General Host Software Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.2.1 GCC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.2 GNU Make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.3 GNU makeinfo Version Requirements . . . . . . . . . . . . . . . . . . . . . . 6

2.3 Host Specific Notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.1 Solaris 2.x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3.2 Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2.4 Archive and Build Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.4.1 RPM Archive and Build Directory Format . . . . . . . . . . . . . . . . . 7 2.4.2 Archive and Build Directory Format . . . . . . . . . . . . . . . . . . . . . . . 7

3 Prebuilt Toolset Executables . . . . . . . . . . . . . . . . . . 9

3.1 RPMs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.1 Installing RPMs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.2 Determining Which RTEMS RPMs are Installed . . . . . . . . . . 10 3.1.3 Removing RPMs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.2 Zipped Tar Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2.1 Installing Zipped Tar Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2.2 Removing Zipped Tar Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Building the GNU Cross Compiler Toolset

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4.1 Building BINUTILS GCC and NEWLIB . . . . . . . . . . . . . . . . . . . . . . 13

4.1.1 Obtain Source and Patches for BINUTILS GCC and

NEWLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.1.2 Unarchiving the Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.1.3 Applying RTEMS Patches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.1.4 Compiling and Installing BINUTILS GCC and NEWLIB . . 16

4.1.4.1 Using RPM to Build BINUTILS GCC and NEWLIB

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.1.4.2 Using configure and make . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.2 Building the GNU Debugger GDB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

4.2.1 Obtain Source and Patches for GDB. . . . . . . . . . . . . . . . . . . . . . 21 4.2.2 Unarchiving the GDB Distribution . . . . . . . . . . . . . . . . . . . . . . . 21 4.2.3 Applying RTEMS Patch to GDB . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2.4 Compiling and Installing the GNU Debugger GDB. . . . . . . . 21

4.2.4.1 Using RPM to Build GDB. . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.2.4.2 Using the GDB configure Script Directly. . . . . . . . . . . . . 23

4.3 Common Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.3.1 Error Message Indicates Invalid Option to Assembler . . . . . 23 4.3.2 Error Messages Indicating Configuration Problems . . . . . . . . 24

5 Building RTEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

5.1 Obtain the RTEMS Source Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.2 Unarchive the RTEMS Source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.3 Add <INSTALL POINT>/bin to Executable PATH. . . . . . . . . . . . 25 5.4 Verifying the Operation of the Cross Toolset. . . . . . . . . . . . . . . . . . . 25 5.5 Building RTEMS for a Specific Target and BSP . . . . . . . . . . . . . . . 26

5.5.1 Using the RTEMS configure Script Directly. . . . . . . . . . . . . . . 26

6 Building the Sample Applications. . . . . . . . . . . . 29

6.1 Set the Environment Variable RTEMS MAKEFILE PATH …. 29 6.2 Executing the Sample Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.3 C/C++ Sample Applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 6.4 Ada Sample Applications. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.5 Build the Sample Application . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.6 Application Executable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.7 More Information on RTEMS Application Makefiles . . . . . . . . . . . 32

7 Where To Go From Here . . . . . . . . . . . . . . . . . . . . . 33

7.1 Documentation Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.2 Writing an Application . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

Appendix A Using MS-Windows as a

Development Host . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

A.1 Microsoft Windows Version Requirements . . . . . . . . . . . . . . . . . . . . . 35 A.2 Cygwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 A.3 Text Editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 A.4 System Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

1 Introduction

The purpose of this document is to guide you through the process of installing a GNU cross development environment to use with RTEMS.

If you are already familiar with the concepts behind a cross compiler and have a background in Unix, these instructions should provide the bare essentials for performing a setup of the following items:

• GNU Cross Compilation Tools for RTEMS on your build-host system • RTEMS OS for the target

• GDB Debugger

The remainder of this chapter provides background information on real-time embedded systems and cross development and an overview of other resources of interest on the Internet. If you are not familiar with real-time embedded systems or the other areas, please read those sections. These sections will help familiarize you with the types of systems RTEMS is designed to be used in and the cross development process used when developing RTEMS applications.

1.1 Real-Time Embedded Systems

Real-time embedded systems are found in practically every facet of our everyday lives. Today’s systems range from the common telephone, automobile control systems, and kitchen appliances to complex air traffic control systems, military weapon systems, and production line control including robotics and automation. However, in the current climate of rapidly changing technology, it is difficult to reach a consensus on the definition of a real-time embedded system. Hardware costs are continuing to rapidly decline while at the same time the hardware is increasing in power and functionality. As a result, embedded systems that were not considered viable two years ago are suddenly a cost effective solution. In this domain, it is not uncommon for a single hardware configuration to employ a variety of architectures and technologies. Therefore, we shall define an embedded system as any computer system that is built into a larger system consisting of multiple technologies such as digital and analog electronics, mechanical devices, and sensors.

Even as hardware platforms become more powerful, most embedded systems are critically dependent on the real-time software embedded in the systems themselves. Regardless of how efficiently the hardware operates, the performance of the embedded real-time software determines the success of the system. As the complexity of the embedded hardware platform grows, so does the size and complexity of the embedded software. Software systems must routinely perform activities which were only dreamed of a short time ago. These large, complex, real-time embedded applications now commonly contain one million lines of code or more.

Real-time embedded systems have a complex set of characteristics that distinguish them from other software applications. Real-time embedded systems are driven by and must respond to real world events while adhering to rigorous requirements imposed by the environment with which they interact. The correctness of the system depends not only on the results of computations, but also on the time at which the results are produced. The most important and complex characteristic of real-time application systems is that they must receive and respond to a set of external stimuli within rigid and critical time constraints.

A single real-time application can be composed of both soft and hard real-time components. A typical example of a hard real-time system is a nuclear reactor control system that must not only detect failures, but must also respond quickly enough to prevent a meltdown. This application also has soft real-time requirements because it may involve a man-machine interface. Providing an interactive input to the control system is not as critical as setting off an alarm to indicate a failure condition. However, the interactive system component must respond within an acceptable time limit to allow the operator to interact efficiently with the control system.

1.2 Cross Development

Today almost all real-time embedded software systems are developed in a cross development environment using cross development tools. In the cross development environment, software development activities are typically performed on one computer system, the build-host system, while the result of the development effort (produced by the cross tools) is a software system that executes on the target platform. The requirements for the target platform are usually incompatible and quite often in direct conflict with the requirements for the buildhost. Moreover, the target hardware is often custom designed for a particular project. This means that the cross development toolset must allow the developer to customize the tools to address target specific run-time issues. The toolset must have provisions for board dependent initialization code, device drivers, and error handling code.

The build-host computer is optimized to support the code development cycle with support for code editors, compilers, and linkers requiring large disk drives, user development windows, and multiple developer connections. Thus the build-host computer is typically a traditional UNIX workstation such as those available from SUN or Silicon Graphics, or a PC running either a version of MS-Windows or UNIX. The build-host system may also be required to execute office productivity applications to allow the software developer to write documentation, make presentations, or track the project’s progress using a project management tool. This necessitates that the build-host computer be general purpose with resources such as a thirty-two or sixty-four bit processor, large amounts of RAM, a monitor, mouse, keyboard, hard and floppy disk drives, CD-ROM drive, and a graphics card. It is likely that the system will be multimedia capable and have some networking capability.

Conversely, the target platform generally has limited traditional computer resources. The hardware is designed for the particular functionality and requirements of the embedded system and optimized to perform those tasks effectively. Instead of hard drives and keyboards, it is composed of sensors, relays, and stepper motors. The per-unit cost of the target platform is typically a critical concern. No hardware component is included without being cost justified. As a result, the processor of the target system is often from a different processor family than that of the build-host system and usually has lower performance. In addition to the processor families designed only for use in embedded systems, there are versions of nearly every general-purpose processor specifically tailored for real-time embedded systems. For example, many of the processors targeting the embedded market do not include hardware floating point units, but do include peripherals such as timers, serial controllers, or network interfaces.

1.3 Resources on the Internet

This section describes various resources on the Internet which are of use to RTEMS users.

1.3.1 Online Tool Documentation

Each of the tools in the GNU development suite comes with documentation. It is in the reader’s and tool maintainers’ interest that one read the documentation before posting a problem to a mailing list or news group. The RTEMS Project provides formatted documentation for the primary tools in the cross development toolset including BINUTILS, GCC, NEWLIB, and GDB with the pre-built versions of those tools.

Much of the documentation is available at other sites on the Internet. The following is a list of URLs where one can find HTML versions of the GNU manuals:

Free Software Foundation http://www.gnu.org/manual/manual.html

                   (http://www.gnu.org/manual/manual.html)

Delorie  Software  http://www.delorie.com/gnu/docs

                   (http://www.delorie.com/gnu/docs)

1.3.2 RTEMS Mailing List

rtems-users@rtems.com (mailto:rtems-users@rtems.com)

This mailing list is dedicated to the discussion of issues related to RTEMS, including GNAT/RTEMS. If you have questions about RTEMS, wish to make suggestions, or just want to pick up hints, this is a good list to monitor. Subscribe by sending an empty mail mes-

sage to rtems-users-subscribe@rtems.com (mailto:rtems-users-subscribe@rtems.com).

Messages sent to rtems-users@rtems.com (mailto:rtems-users@rtems.com) are posted to the list.

1.3.3 CrossGCC Mailing List

crossgcc@sources.redhat.com (mailto:crossgcc@sources.redhat.com)

This mailing list is dedicated to the use of the GNU tools in cross development environments. Most of the discussions focus on embedded issues. Information on subscribing to this mailing list is included in the CrossGCC FAQ (http://www.objsw.com/CrossGCC/).

The CrossGCC FAQ and Wiki are are available at http://www.billgatliff.com

(http://www.billgatliff.com).

1.3.4 GCC Mailing Lists

The GCC Project is hosted at http://gcc.gnu.org (http://gcc.gnu.org). They maintain multiple mailing lists that are described at the web site along with subscription information.

2 Requirements

This chapter describes the build-host system requirements and initial steps in installing the GNU Cross Compiler Tools and RTEMS on a build-host.

2.1 Disk Space

A fairly large amount of disk space is required to perform the build of the GNU C/C++ Cross Compiler Tools for RTEMS. The following table may help in assessing the amount of disk space required for your installation:

     +------------------------------------+--------------------------+
     |              Component             |   Disk Space Required    |
     +------------------------------------+--------------------------+
     |        archive directory           |        55 Mbytes         |
     |        tools src unarchived        |       350 Mbytes         |
     |  each individual build directory   |     up to 750 Mbytes     |
     |     each installation directory    |      20-200 Mbytes       |
     +------------------------------------+--------------------------+

It is important to understand that the above requirements only address the GNU C/C++ Cross Compiler Tools themselves. Adding additional languages such as Fortran or Objective-C can increase the size of the build and installation directories. Also, the unarchived source and build directories can be removed after the tools are installed.

After the tools themselves are installed, RTEMS must be built and installed for each Board Support Package that you wish to use. Thus the precise amount of disk space required for each installation directory depends highly on the number of RTEMS BSPs which are to be installed. If a single BSP is installed, then the additional size of each install directory will tend to be in the 40-60 Mbyte range.

There are a number of factors which must be taken into account in order to estimate the amount of disk space required to build RTEMS itself. Attempting to build multiple BSPs in a single step increases the disk space requirements. Similarly enabling optional features increases the build and install space requirements. In particular, enabling and building the RTEMS tests results in a significant increase in build space requirements but since the tests are not installed has, enabling them has no impact on installation requirements.

2.2 General Host Software Requirements

The instructions in this manual should work on any computer running a UNIX variant. Some native GNU tools are used by this procedure including:

• GCC

• GNU make

• GNU makeinfo

In addition, some native utilities may be deficient for building the GNU tools.

2.2.1 GCC

Although RTEMS itself is intended to execute on an embedded target, there is source code for some native programs included with the RTEMS distribution. Some of these programs are used to assist in the building of RTEMS itself, while others are BSP specific tools. Regardless, no attempt has been made to compile these programs with a non-GNU compiler.

2.2.2 GNU Make

Both NEWLIB and RTEMS use GNU make specific features and can only be built using GNU make. Many systems include a make utility that is not GNU make. The safest way to meet this requirement is to ensure that when you invoke the command make, it is GNU make. This can be verified by attempting to print the GNU make version information:

     make --version

If you have GNU make and another make on your system, it is common to put the directory containing GNU make before the directory containing other implementations of make.

2.2.3 GNU makeinfo Version Requirements

In order to build gcc 2.9.x or newer versions, the GNU makeinfo program installed on your system must be at least version 1.68. The appropriate version of makeinfo is distributed with gcc.

The following demonstrates how to determine the version of makeinfo on your machine:

     makeinfo --version

2.3 Host Specific Notes

2.3.1 Solaris 2.x

The following problems have been reported by Solaris 2.x users:

• The build scripts are written in "shell". The program /bin/sh on Solaris 2.x is not robust enough to execute these scripts. If you are on a Solaris 2.x host, then use the /bin/ksh or /bin/bash shell instead.

• The native patch program is broken. Install the GNU version.

• The native m4 program is deficient. Install the GNU version.

2.3.2 Linux

The following problems have been reported by Linux users:

• Certain versions of GNU fileutils include a version of install which does not work

properly. Please perform the following test to see if you need to upgrade:

           install -c -d /tmp/foo/bar

If this does not create the specified directories your install program will not install RTEMS properly. You will need to upgrade to at least GNU fileutils version 3.16 to resolve this problem.

2.4 Archive and Build Directories

If you are using RPM or another packaging format that supports building a package from source, then there is probably a directory structure assumed by that packaging format. Otherwise, you are free to use whatever organization you like. However, this document will use the directory organization described in Section 2.4.2 [Archive and Build Directory Format], page 7.

2.4.1 RPM Archive and Build Directory Format

For RPM, it is assumed that the following subdirectories are under a root directory such as

/usr/src/redhat:
     BUILD
     RPMS
     SOURCES
     SPECS
     SRPMS

For the purposes of this document, the RPM SOURCES directory is the directory into which all tool source and patches are assumed to reside. The BUILD directory is where the actual build is performed when building binaries from a source RPM. The SOURCES and BUILD are logically equivalent to the archive and tools directory discussed in the next section.

2.4.2 Archive and Build Directory Format

When no packaging format requirements are present, the root directory for the storage of source archives and patches as well as for building the tools is up to the user. The only concern is that there be enough disk space to complete the build. In this document, the following organization will be used.

Make an archive directory to contain the downloaded source code and a tools directory to be used as a build directory. The command sequence to do this is shown below:

     mkdir archive
     mkdir tools

This will result in an initial directory structure similar to the one shown in the following figure:

     /whatever/prefix/you/choose/
             archive/
             tools/

3 Prebuilt Toolset Executables

Precompiled toolsets are available for Linux, Cygwin, FreeBSD, and Solaris. These are packaged in the following formats:

• Linux - RPM

• Cygwin - tar.bz2

• Solaris - tar.bz2

RPM is an acronym for the RPM Package Manager. RPM is the native package installer for many Linux distributions including RedHat, SuSE, and Fedora.

The prebuilt binaries are intended to be easy to install and the instructions are similar regardless of the host environment. There are a few structural issues with the packaging of the RTEMS Cross Toolset binaries that you need to be aware of.

1. There are dependencies between the various packages. This requires that certain packages be installed before others may be. Some packaging formats enforce this dependency.

2. Some packages are target CPU family independent and shared across all target architectures. These are referred to as "base" packages.

3. Depending upon the version of GCC as well as the development host and target CPU combination, pre-built supplemental packages may be provided for Ada (gnat), Chill, Java (gcj), Fortran (g77), and Objective-C (objc). These binaries are strictly optional.

NOTE: Installing toolset binaries does not install RTEMS itself, only the tools required to build RTEMS. See Chapter 5 [Building RTEMS], page 25 for the next step in the process.

3.1 RPMs

This section provides information on installing and removing RPMs.

3.1.1 Installing RPMs

The following is a sample session illustrating the installation of a C/C++ toolset targeting the SPARC architecture.

     rpm -U rtems-4.9-binutils-common-2.18-4.i386.rpm
     rpm -U rtems-4.9-sparc-rtems4.9-binutils-2.18-4.i386.rpm
     rpm -U rtems-4.9-gcc-common-4.3.2-15.i386.rpm
     rpm -U rtems-4.9-sparc-rtems4.9-newlib-1.16.0-15.i386.rpm
     rpm -U rtems-4.9-sparc-rtems4.9-gcc-4.3.2-15.i386.rpm
     rpm -U rtems-4.9-sparc-rtems4.9-gcc-c++-4.3.2-15.i386.rpm
     rpm -U rtems-4.9-gdb-common-6.8-4.i386.rpm
     rpm -U rtems-4.9-sparc-rtems4.9-gdb-6.8-4.i386.rpm

Upon successful completion of the above command sequence, a C/C++ cross development toolset targeting the SPARC is installed in /opt/rtems-4.9.1. In order to use this toolset, the directory /opt/rtems-4.9.1/bin must be included in your PATH.

Once you have successfully installed the RPMs for BINUTILS, GCC, NEWLIB, and GDB, then you may proceed directly to Chapter 5 [Building RTEMS], page 25.

3.1.2 Determining Which RTEMS RPMs are Installed

The following command will report which RTEMS RPMs are currently installed:

rpm -q -g RTEMS/4.9

3.1.3 Removing RPMs

The following is a sample session illustrating the removal of a C/C++ toolset targeting the SPARC architecture.

rpm -e rtems-4.9-sparc-rtems4.9-gdb
rpm -e rtems-4.9-gdb-common
rpm -e rtems-4.9-sparc-rtems4.9-gcc-c++
rpm -e rtems-4.9-sparc-rtems4.9-gcc
rpm -e rtems-4.9-sparc-rtems4.9-newlib
rpm -e rtems-4.9-gcc-common
rpm -e rtems-4.9-sparc-rtems4.9-binutils
rpm -e rtems-4.9-binutils-common

NOTE: If you have installed any RTEMS BSPs, then it is likely that RPM will complain about not being able to remove everything.

3.2 Zipped Tar Files

This section provides information on installing and removing Zipped Tar Files (e.g .tar.gz or .tar.bz2).

3.2.1 Installing Zipped Tar Files

The following is a sample session illustrating the installation of a C/C++ toolset targeting the SPARC architecture assuming that GNU tar is installed as tar for a set of archive files compressed with GNU Zip (gzip):

cd /
tar xzf rtems-4.9-binutils-common-2.18-4.tar.gz
tar xzf rtems-4.9-sparc-rtems4.9-binutils-2.18-4.tar.gz
tar xzf rtems-4.9-gcc-common-4.3.2-15.tar.gz
tar xzf rtems-4.9-sparc-rtems4.9-gcc-4.3.2-15.tar.gz
tar xzf rtems-4.9-sparc-rtems4.9-newlib-1.16.0-15.tar.gz
tar xzf rtems-4.9-gdb-common-6.8-4.tar.gz
tar xzf rtems-4.9-sparc-rtems4.9-gdb-6.8-4.tar.gz

The following command set is the equivalent command sequence for the same toolset assuming that is was compressed with GNU BZip (bzip2):

cd /
tar xjf rtems-4.9-binutils-common-2.18-4.tar.bz2
tar xjf rtems-4.9-sparc-rtems4.9-binutils-2.18-4.tar.bz2
tar xjf rtems-4.9-gcc-common-4.3.2-15.tar.bz2

     tar xjf rtems-4.9-sparc-rtems4.9-newlib-1.16.0-15.tar.bz2
     tar xjf rtems-4.9-sparc-rtems4.9-gcc-4.3.2-15.tar.bz2
     tar xjf rtems-4.9-gdb-common-6.8-4.tar.bz2
     tar xjf rtems-4.9-sparc-rtems4.9-gdb-6.8-4.tar.bz2

Upon successful completion of the above command sequence, a C/C++ cross development toolset targeting the SPARC is installed in /opt/rtems-4.9.1. In order to use this toolset, the directory /opt/rtems-4.9.1 must be included in your PATH.

3.2.2 Removing Zipped Tar Files

There is no automatic way to remove the contents of a tar.gz or tar.bz2 once it is installed. The contents of the directory /opt/rtems-4.9.1 can be removed but this will likely result in other packages being removed as well.

4 Building the GNU Cross Compiler Toolset

NOTE: This chapter does NOT apply if you installed prebuilt toolset executables for BINU-TILS, GCC, NEWLIB, and GDB. If you installed prebuilt executables for all of those, proceed to Chapter 5 [Building RTEMS], page 25. If you require a GDB with a special configuration to connect to your target board, then proceed to Section 4.2 [Building the GNU Debugger GDB], page 20 for some advice.

This chapter describes the steps required to acquire the source code for a GNU cross compiler toolset, apply any required RTEMS specific patches, compile that toolset and install it.

It is recommended that when toolset binaries are available for your particular host, that they be used. Prebuilt binaries are much easier to install.

4.1 Building BINUTILS GCC and NEWLIB

NOTE: This step is NOT required if prebuilt executables for BINUTILS, GCC, and NEWLIB were installed.

This section describes the process of building BINUTILS, GCC, and NEWLIB using a variety of methods. Included is information on obtaining the source code and patches, applying patches, and building and installing the tools using multiple methods.

4.1.1 Obtain Source and Patches for BINUTILS GCC and NEWLIB

NOTE: This step is required for all methods of building BINUTILS, GCC, and NEWLIB.

This section lists the components required to build BINUTILS, GCC, and NEWLIB from source to target RTEMS. These files should be placed in your archive directory. Included are the locations of each component as well as any required RTEMS specific patches.

gcc-4.3.2

         FTP Site:    ftp.gnu.org
         Directory:   /pub/gnu/gcc/gcc-4.3.2
         File:        gcc-4.3.2.tar.bz2
         URL:         ftp://ftp.gnu.org/pub/gnu/gcc/gcc-4.3.2/gcc-4.3.2.tar.bz2

binutils-2.18

         FTP Site:    ftp.gnu.org
         Directory:   /pub/gnu/binutils
         File:        binutils-2.18.tar.gz
         URL:         ftp://ftp.gnu.org/pub/gnu/binutils/binutils-2.18.tar.gz

newlib-1.16.0

         FTP Site:    sources.redhat.com
         Directory:   /pub/newlib
         File:        newlib-1.16.0.tar.gz
         URL:         ftp://sources.redhat.com/pub/newlib/newlib-1.16.0.tar.gz

RTEMS Specific Tool Patches and Scripts

         FTP Site:    ftp.rtems.com
         Directory:   /pub/rtems/SOURCES/4.9
         File:        binutils-2.18-rtems4.9-20080211.diff
         File:        newlib-1.16.0-rtems4.9-20080827.diff
         File:        gcc-4.3.2-rtems4.9-20080917.diff

The individual components in the RTEMS Development Toolset are updated independently of one another by their respective maintainers. In addition, the patches which the RTEMS Project has determined are necessary are subject to change as users report issues on individual host and target platforms. As a result, it is possible that the versions listed in this document are NOT the latest ones available. The latest patches for each tool are always available from RTEMS CVS on the respective branch and should always be mirrored on the RTEMS ftp site. It is recommended that before building a toolset from source, you verify you are using the latest patches.

4.1.2 Unarchiving the Tools

NOTE: This step is required if building BINUTILS, GCC, and NEWLIB using the procedure described in Section 4.1.4.2 [Using configure and make], page 18. It is NOT required if using the procedure described in Section 4.1.4.1 [Using RPM to Build BINUTILS GCC and NEWLIB], page 16.

GNU source distributions are archived using tar and compressed using either gzip or bzip. If compressed with gzip, the extension .gz is used. If compressed with bzip, the extension .bz2 is used.

While in the tools directory, unpack the compressed tar files for BINUTILS, GCC, and NEWLIB using the appropriate command based upon the compression program used.

     cd tools
     tar xzf ../archive/TOOLNAME.tar.gz  # for gzip'ed tools
     tar xjf ../archive/TOOLNAME.tar.bz2 # for bzip'ed tools

After the compressed tar files have been unpacked using the appropriate commands, the following directories will have been created under tools.

• binutils-2.18

• gcc-4.3.2

• newlib-1.16.0

The tree should look something like the following figure:

     /whatever/prefix/you/choose/
             archive/
                 gcc-4.3.2.tar.bz2
                 binutils-2.18.tar.gz
                 newlib-1.16.0.tar.gz
                 gcc-4.3.2-rtems4.9-20080917.diff
                 binutils-2.18-rtems4.9-20080211.diff
                 newlib-1.16.0-rtems4.9-20080827.diff
             tools/
                 binutils-2.18/
                 gcc-4.3.2/
                 newlib-1.16.0/

4.1.3 Applying RTEMS Patches

NOTE: This step is required if building BINUTILS, GCC, and NEWLIB using the procedures described in Section 4.1.4.2 [Using configure and make], page 18. It is NOT required if using the procedure described in Section 4.1.4.1 [Using RPM to Build BINUTILS GCC and NEWLIB], page 16.

This section describes the process of applying the RTEMS patches to GCC, NEWLIB, and BINUTILS.

Apply RTEMS Patch to GCC

Apply the patch using the following command sequence:

     cd tools/gcc-4.3.2
     cat ../../archive/gcc-4.3.2-rtems4.9-20080917.diff | \
         patch -p1

If the patch was compressed with the gzip program, it will have a suffix of .gz and you should use zcat instead of cat as shown above. If the patch was compressed with the gzip program, it will have a suffix of .bz2 and you should use bzcat instead of cat as shown above.

Check to see if any of these patches have been rejected using the following sequence:

     cd tools/gcc-4.3.2
     find . -name "*.rej" -print

If any files are found with the .rej extension, a patch has been rejected. This should not happen with a good patch file which is properly applied.

Apply RTEMS Patch to binutils

Apply the patch using the following command sequence:

     cd tools/binutils-2.18
     cat ../../archive/binutils-2.18-rtems4.9-20080211.diff | \
         patch -p1

If the patch was compressed with the gzip program, it will have a suffix of .gz and you should use zcat instead of cat as shown above. If the patch was compressed with the gzip program, it will have a suffix of .bz2 and you should use bzcat instead of cat as shown above.

Check to see if any of these patches have been rejected using the following sequence:

     cd tools/binutils-2.18
     find . -name "*.rej" -print

If any files are found with the .rej extension, a patch has been rejected. This should not happen with a good patch file which is properly applied.

Apply RTEMS Patch to newlib

Apply the patch using the following command sequence:

     cd tools/newlib-1.16.0
     cat ../../archive/newlib-1.16.0-rtems4.9-20080827.diff | \
         patch -p1

If the patch was compressed with the gzip program, it will have a suffix of .gz and you should use zcat instead of cat as shown above. If the patch was compressed with the gzip program, it will have a suffix of .bz2 and you should use bzcat instead of cat as shown above.

Check to see if any of these patches have been rejected using the following sequence:

     cd tools/newlib-1.16.0
     find . -name "*.rej" -print

If any files are found with the .rej extension, a patch has been rejected. This should not happen with a good patch file which is properly applied.

4.1.4 Compiling and Installing BINUTILS GCC and NEWLIB

There are two supported methods to compile and install BINUTILS, GCC, and NEWLIB:

• RPM

• direct invocation of configure and make

Direct invocation of configure and make provides more control and easier recovery from problems when building.

4.1.4.1 Using RPM to Build BINUTILS GCC and NEWLIB

NOTE: The procedures described in the following sections must be completed before this step:

• Section 4.1.1 [Obtain Source and Patches for BINUTILS GCC and NEWLIB], page 13

RPM automatically unarchives the source and applies any needed patches so you do NOT have to manually perform the procedures described Section 4.1.2 [Unarchiving the Tools], page 14 and Section 4.1.3 [Applying RTEMS Patches], page 15.

This section describes the process of building binutils, gcc, and newlib using RPM. RPM is a packaging format which can be used to distribute binary files as well as to capture the procedure and source code used to produce those binary files. Before attempting to build any RPM from source, it is necessary to ensure that all required source and patches are in the SOURCES directory under the RPM root (probably /usr/src/redhat or /usr/local/src/redhat) on your machine. This procedure starts by installing the source RPMs as shown in the following example:

     rpm -U rtems-4.9-i386-rtems4.9-binutils-2.18-4.src.rpm
     rpm -U rtems-4.9-i386-rtems4.9-gcc-4.3.2-15.src.rpm

The RTEMS tool source RPMS are called "nosrc" to indicate that one or more source files required to produce the RPMs are not present. The RTEMS source RPMs typically include all required patches, but do not include the large .tar.gz or .tgz files for each component such as BINUTILS, GCC, or NEWLIB. These are shared by all RTEMS RPMs regardless of target CPU and there was no reason to duplicate them. You will have to get the required source archive files by hand and place them in the SOURCES directory before attempting to build. If you forget to do this, RPM is smart – it will tell you what is missing. To determine what is included or referenced by a particular RPM, use a command like the following:

     $ rpm -q -l -p rtems-4.9-i386-rtems4.9-gcc-3.2.3-1.src.rpm
     gcc-3.2.3-rtems4.9-20030507a.diff
     rtems-4.9-i386-rtems4.7-gcc.spec
     newlib-1.11.0-rtems4.9-20030507.diff

Notice that there are patch files (the .diff files) and a file describing the build procedure and files produced (the .spec file), but no source archives (the *tar.*

files). When installing this source RPM (rpm -U rtems-4.9-i386-rtems4.9-gcc-newlib-
gcc3.2.3newlib1.11.0-1.src.rpm), the .spec file is placed in the SPECS directory under

the RPM root directory, while the *.diff files are placed into the SOURCES directory.

Configuring and Building BINUTILS using RPM

The following example illustrates the invocation of RPM to build a new, locally compiled, binutils binary RPM that matches the installed source RPM. This example assumes that all of the required source is installed.

     cd <RPM_ROOT_DIRECTORY>/SPECS
     rpm -bb i386-rtems4.9-binutils-2.18.spec

If the build completes successfully, RPMS like the following will be generated in a build-host architecture specific subdirectory of the RPMS directory under the RPM root directory.

     rtems-4.9-binutils-common-2.18-4.i386.rpm
     rtems-4.9-i386-rtems4.9-binutils-2.18-4.i386.rpm

NOTE: It may be necessary to remove the build tree in the BUILD directory under the RPM root directory.

Configuring and Building GCC and NEWLIB using RPM

The following example illustrates the invocation of RPM to build a new, locally compiled, set of GCC and NEWLIB binary RPMs that match the installed source RPM. It is also necessary to install the BINUTILS RPMs and place them in your PATH. This example assumes that all of the required source is installed.

     cd <RPM_ROOT_DIRECTORY>/RPMS/i386
     rpm -U rtems-4.9-binutils-common-2.18-4.i386.rpm
     rpm -U rtems-4.9-i386-rtems4.9-binutils-2.18-4.i386.rpm
     export PATH=/opt/rtems-4.9.1/bin:$PATH
     cd <RPM_ROOT_DIRECTORY>/SPECS
     rpm -bb i386-rtems4.9-gcc-4.3.2-newlib-1.16.0.spec

If the build completes successfully, a set of RPMS like the following will be generated in a build-host architecture specific subdirectory of the RPMS directory under the RPM root directory.

     rtems-4.9-gcc-common-4.3.2-15.i386.rpm
     rtems-4.9-i386-rtems4.9-newlib-1.16.0-15.i386.rpm
     rtems-4.9-i386-rtems4.9-gcc-4.3.2-15.i386.rpm
     rtems-4.9-i386-rtems4.9-gcc-c++-4.3.2-15.i386.rpm

NOTE: Some targets do not support building all languages.

NOTE: It may be necessary to remove the build tree in the BUILD directory under the RPM root directory.

4.1.4.2 Using configure and make

NOTE: The procedures described in the following sections must be completed before this step:

• Section 4.1.1 [Obtain Source and Patches for BINUTILS GCC and NEWLIB], page 13

• Section 4.1.2 [Unarchiving the Tools], page 14

• Section 4.1.3 [Applying RTEMS Patches], page 15

This section describes the process of building binutils, gcc, and newlib manually using

configure and make directly.

Configuring and Building BINUTILS

The following example illustrates the invocation of configure and make to build and install binutils-2.18 for the sparc-rtems4.9 target:

     mkdir b-binutils
     cd b-binutils
     ../binutils-2.18/configure --target=sparc-rtems4.9 \
       --prefix=/opt/rtems-4.9.1
     make all
     make info
     make install

After binutils-2.18 is built and installed the build directory b-binutils may be removed.

For more information on the invocation of configure, please refer to the documentation for binutils-2.18 or invoke the binutils-2.18 configure command with the --help option.

NOTE: The shell PATH variable needs to be updated to include the path the binutils user executables have been installed in. The directory containing the executables is the prefix used above with ‘bin’ post-fixed.

     export PATH=/opt/rtems-4.9.1/bin:${PATH}

Failure to have the binutils in the path will cause the GCC and NEWLIB build to fail with an error message similar to:

     sparc-rtems4.9-ar: command not found

Configuring and Building GCC and NEWLIB

Before building gcc-4.3.2 and newlib-1.16.0, binutils-2.18 must be installed and the directory containing those executables must be in your PATH.

The C Library is built as a subordinate component of gcc-4.3.2. Because of this, the newlib-1.16.0 directory source must be available inside the gcc-4.3.2 source tree. This is normally accomplished using a symbolic link as shown in this example:

     cd gcc-4.3.2
     ln -s ../newlib-1.16.0/newlib .

The following example illustrates the invocation of configure and make to build and install gcc-4.3.2 with only C and C++ support for the sparc-rtems4.9 target:

     mkdir b-gcc
     cd b-gcc
     ../gcc-4.3.2/configure --target=sparc-rtems4.9 \
        --with-gnu-as --with-gnu-ld --with-newlib --verbose \
        --enable-threads --enable-languages="c,c++" \
        --prefix=/opt/rtems-4.9.1
     make all
     make info
     make install

After gcc-4.3.2 is built and installed the build directory b-gcc may be removed.

For more information on the invocation of configure, please refer to the documentation for gcc-4.3.2 or invoke the gcc-4.3.2 configure command with the --help option.

Building GCC with Ada Support

If you want a GCC toolset that includes support for Ada (e.g. GNAT), there are some additional requirements on the host environment and additional build steps to perform. It is critical that you use the same version of GCC/GNAT as the native compiler. GNAT must be compiled with an Ada compiler and when building a GNAT cross-compiler, it should be the same version of GNAT itself.

The build procedure is the same until the configure step. A GCC toolset with GNAT enabled requires that ada be included in the set of enabled languages. The following example illustrates the invocation of configure and make to build and install gcc-4.3.2 with only C, C++, and Ada support for the sparc-rtems4.9 target:

     mkdir b-gcc
     cd gcc-4.3.2/gcc/ada
     touch treeprs.ads [es]info.h nmake.ad[bs]
     cd ../../../b-gcc
     ../gcc-4.3.2/configure --target=sparc-rtems4.9 \
        --with-gnu-as --with-gnu-ld --with-newlib --verbose \
        --enable-threads --enable-languages="c,c++,ada" \
        --prefix=/opt/rtems-4.9.1
     make all
     make info
     make -C gcc cross-gnattools
     make -C gcc ada.all.cross
     make -C gcc GNATLIBCFLAGS="USER_SELECTED_CPU_CFLAGS" gnatlib
     make install

After gcc-4.3.2 is built and installed the build directory b-gcc may be removed.

4.2 Building the GNU Debugger GDB

NOTE: This step is NOT required if prebuilt executables for the GNU Debugger GDB were installed.

The GNU Debugger GDB supports many configurations but requires some means of communicating between the host computer and target board. This communication can be via a serial port, Ethernet, BDM, or ROM emulator. The communication protocol can be the GDB remote protocol or GDB can talk directly to a ROM monitor. This setup is target board specific. The following configurations have been successfully used with RTEMS applications:

• BDM with ColdFire, 683xx, MPC860 CPUs

• Motorola Mxxxbug found on M68xxx VME boards

• Motorola PPCbug found on PowerPC VME, CompactPCI, and MTX boards

• ARM based Cogent EDP7312

• PC’s using various Intel and AMD CPUs including i386, i486, Pentium and above, and Athlon

• PowerPC Instruction Simulator in GDB (PSIM)

• MIPS Instruction Simulator in GDB (JMR3904)

• Sparc Instruction Simulator in GDB (SIS)

• Sparc Instruction Simulator (TSIM)

• DINK32 on various PowerPC boards

GDB is currently RTEMS thread/task aware only if you are using the remote debugging support via Ethernet. These are configured using gdb targets of the form CPU-RTEMS. Note the capital RTEMS.

It is recommended that when toolset binaries are available for your particular host, that they be used. Prebuilt binaries are much easier to install but in the case of gdb may or may not include support for your particular target board.

4.2.1 Obtain Source and Patches for GDB

NOTE: This step is required for all methods of building GDB.

This section lists the components required to build GDB from source to target RTEMS. These files should be placed in your archive directory. Included are the locations of each component as well as any required RTEMS specific patches.

gdb-6.8

         FTP Site:    ftp.gnu.org
         Directory:   /pub/gnu/gdb
         File:        gdb-6.8.tar.gz
         URL:         ftp://ftp.gnu.org/pub/gnu/gdb/gdb-6.8.tar.gz

RTEMS Specific Tool Patches and Scripts

         FTP Site:    ftp.rtems.com
         Directory:   /pub/rtems/SOURCES/4.9
         File:        gdb-6.8-rtems4.9-20080917.diff
         URL:         ftp://ftp.rtems.com/pub/rtems/SOURCES/gdb-6.8-rtems4.9-20080917.diff

4.2.2 Unarchiving the GDB Distribution

Use the following commands to unarchive the GDB distribution:

     cd tools
     tar xzf ../archive/gdb-6.8.tar.gz

The directory gdb-6.8 is created under the tools directory.

4.2.3 Applying RTEMS Patch to GDB

Apply the patch using the following command sequence:

     cd tools/gdb-6.8
     cat archive/gdb-6.8-rtems4.9-20080917.diff | \
         patch -p1

If the patch was compressed with the gzip program, it will have a suffix of .gz and you should use zcat instead of cat as shown above. If the patch was compressed with the gzip program, it will have a suffix of .bz2 and you should use bzcat instead of cat as shown above.

Check to see if any of these patches have been rejected using the following sequence:

     cd tools/gdb-6.8
     find . -name "*.rej" -print

If any files are found with the .rej extension, a patch has been rejected. This should not happen with a good patch file.

4.2.4 Compiling and Installing the GNU Debugger GDB

There are three methods of building the GNU Debugger:

• RPM

• direct invocation of configure and make

Direct invocation of configure and make provides more control and easier recovery from problems when building.

4.2.4.1 Using RPM to Build GDB

This section describes the process of building binutils, gcc, and newlib using RPM. RPM is a packaging format which can be used to distribute binary files as well as to capture the procedure and source code used to produce those binary files. Before attempting to build any RPM from source, it is necessary to ensure that all required source and patches are in the SOURCES directory under the RPM root (probably /usr/src/redhat or /usr/local/src/redhat) on your machine. This procedure starts by installing the source RPMs as shown in the following example:

     rpm -U rtems-4.9-i386-rtems4.9-gdb-6.8-4.src.rpm

Because RTEMS tool RPMS are called "nosrc" to indicate that one or more source files required to produce the RPMs are not present. The RTEMS source GDB RPM does not include the large .tar.gz or .tgz files for GDB. This is shared by all RTEMS RPMs regardless of target CPU and there was no reason to duplicate them. You will have to get the required source archive files by hand and place them in the SOURCES directory before attempting to build. If you forget to do this, RPM is smart – it will tell you what is missing. To determine what is included or referenced by a particular RPM, use a command like the following:

     $ rpm -q -l -p rtems-4.9-i386-rtems4.9-gdb-6.8-4.src.rpm
     gdb-6.8-rtems4.9-rtems4.9-20080917.diff
     gdb-6.8.tar.gz
     i386-rtems4.9-gdb-6.8.spec

Notice that there is a patch file (the .diff file), a source archive file (the .tar.gz), and a file describing the build procedure and files produced (the .spec file). The .spec file is placed in the SPECS directory under the RPM root directory.

Configuring and Building GDB using RPM

The following example illustrates the invocation of RPM to build a new, locally compiled, binutils binary RPM that matches the installed source RPM. This example assumes that all of the required source is installed.

     cd <RPM_ROOT_DIRECTORY>/SPECS
     rpm -bb i386-rtems4.9-gdb-6.8.spec

If the build completes successfully, RPMS like the following will be generated in a build-host architecture specific subdirectory of the RPMS directory under the RPM root directory.

     rtems-4.9-gdb-common-6.8-4.i386.rpm
     rtems-4.9-i386-rtems4.9-gdb-6.8-4.i386.rpm

NOTE: It may be necessary to remove the build tree in the BUILD directory under the RPM root directory.

4.2.4.2 Using the GDB configure Script Directly

This section describes how to configure the GNU debugger for RTEMS targets using configure and make directly. The following example illustrates the invocation of configure and make to build and install gdb-6.8 for the m68k-rtems4.9 target:

     mkdir b-gdb
     cd b-gdb
     ../gdb-6.8/configure --target=m68k-rtems4.9 \
       --prefix=/opt/rtems-4.9.1
     make all
     make info
     make install

For some configurations, it is necessary to specify extra options to configure to enable and configure option components such as a processor simulator. The following is a list of configurations for which there are extra options:

powerpc-rtems4.9  --enable-sim --enable-sim-powerpc --enable-sim-timebase
                   --enable-sim-hardware

sparc-rtems4.9 --enable-sim

After gdb-6.8 is built and installed the build directory b-gdb may be removed.

For more information on the invocation of configure, please refer to the documentation for gdb-6.8 or invoke the gdb-6.8 configure command with the --help option.

4.3 Common Problems

4.3.1 Error Message Indicates Invalid Option to Assembler

If a message like this is printed then the new cross compiler is most likely using the native assembler instead of the cross assembler or vice-versa (native compiler using new cross assembler). This can occur for one of the following reasons:

• Binutils Patch Improperly Applied

• Binutils Not Built

• Current Directory is in Your PATH

If you are using binutils 2.9.1 or newer with certain older versions of gcc, they do not agree on what the name of the newly generated cross assembler is. Older binutils called it as.new which became as.new.exe under Windows. This is not a valid file name, so as.new is now called as-new. By using the latest released tool versions and RTEMS patches, this problem will be avoided.

If binutils did not successfully build the cross assembler, then the new cross gcc (xgcc) used to build the libraries can not find it. Make sure the build of the binutils succeeded.

If you include the current directory in your PATH, then there is a chance that the native compiler will accidentally use the new cross assembler instead of the native one. This usually indicates that "." is before the standard system directories in your PATH. As a general rule, including "." in your PATH is a security risk and should be avoided. Remove "." from your PATH.

NOTE: In some environments, it may be difficult to remove "." completely from your PATH. In this case, make sure that "." is after the system directories containing "as" and "ld".

4.3.2 Error Messages Indicating Configuration Problems

If you see error messages like the following,

• cannot configure libiberty

• coff-emulation not found

• etc.

Then it is likely that one or more of your gnu tools is already configured locally in its source tree. You can check for this by searching for the config.status file in the various tool source trees. The following command does this for the binutils source:

     find binutils-2.18 -name config.status -print

The solution for this is to execute the command make distclean in each of the GNU tools root source directory. This should remove all generated files including Makefiles.

This situation usually occurs when you have previously built the tool source for some non-RTEMS target. The generated configuration specific files are still in the source tree and the include path specified during the RTEMS build accidentally picks up the previous configuration. The include path used is something like this:

     -I../../binutils-2.18/gcc -I/binutils-2.18/gcc/include -I.

Note that the tool source directory is searched before the build directory.

This situation can be avoided entirely by never using the source tree as the build directory – even for

5 Building RTEMS

5.1 Obtain the RTEMS Source Code

This section provides pointers to the RTEMS source code and Hello World example program. These files should be placed in your archive directory.

RTEMS 4.9.1

         FTP Site:    ftp.rtems.com
         Directory:   /pub/rtems/4.9.1
         File:        rtems-4.9.1.tar.bz2
         URL:         ftp://ftp.rtems.com/pub/rtems/4.9.1/rtems-4.9.1.tar.bz2

RTEMS Examples including Hello World

         FTP Site:    ftp.rtems.com
         Directory:   /pub/rtems/4.9.1
         File:        examples-4.9.1.tar.bz2
         URL:         ftp://ftp.rtems.com/pub/rtems/4.9.1/examples-4.9.1.tar.bz2

5.2 Unarchive the RTEMS Source

Use the following command sequence to unpack the RTEMS source into the tools directory:

     cd tools
     tar xjf ../archive/rtems-4.9.1.tar.bz2

This creates the directory rtems-4.9.1.

5.3 Add <INSTALL POINT>/bin to Executable PATH

In order to compile RTEMS, you must have the cross compilation toolset in your search path. It is important to have the RTEMS toolset first in your path to ensure that you are using the intended version of all tools. The following command prepends the directory where the tools were installed in a previous step:

     export PATH=<INSTALL_POINT>/bin:${PATH}

NOTE: The above command is in Bourne shell (sh) syntax and should work with the Korn (ksh) and GNU Bourne Again Shell (bash). It will not work with the C Shell (csh) or derivatives of the C Shell.

5.4 Verifying the Operation of the Cross Toolset

In order to insure that the cross-compiler is invoking the correct subprograms (like as and ld), one can test assemble a small program. When in verbose mode, gcc prints out information showing where it found the subprograms it invokes. In a temporary working directory, place the following function in a file named f.c:

     int f( int x )
     {

       return x + 1;
     }

Then assemble the file using a command similar to the following:

     m68k-rtems-gcc -v -S f.c

Where m68k-rtems-gcc should be changed to match the installed name of your cross compiler. The result of this command will be a sequence of output showing where the crosscompiler searched for and found its subcomponents. Verify that these paths correspond to your <INSTALL POINT>.

Look at the created file f.s and verify that it is in fact for your target processor.

Then try to compile the file f.c directly to object code using a command like the following:

     m68k-rtems-gcc -v -c f.c

If this produces messages that indicate the assembly code is not valid, then it is likely that you have fallen victim to one of the problems described in Section 4.3.1 [Error Message Indicates Invalid Option to Assembler], page 23 Don’t feel bad about this, one of the most common installation errors is for the cross-compiler not to be able to find the cross assembler and default to using the native as. This can result in very confusing error messages.

5.5 Building RTEMS for a Specific Target and BSP

This section describes how to configure and build RTEMS so that it is specifically tailored for your BSP and the CPU model it uses. There is currently only one supported method to compile and install RTEMS:

• direct invocation of configure and make

Direct invocation of configure and make provides more control and easier recovery from problems when building.

This section describes how to build RTEMS.

5.5.1 Using the RTEMS configure Script Directly

Make a build directory under tools and build the RTEMS product in this directory. The ../rtems-4.9.1/configure command has numerous command line arguments. These arguments are discussed in detail in documentation that comes with the RTEMS distribution. A full list of these arguments can be obtained by running ../rtems-4.9.1/configure - -help If you followed the procedure described in the section Section 5.2 [Unarchive the RTEMS Source], page 25, these configuration options can be found in the file tools/rtems-4.9.1/README.configure.

NOTE: The GNAT/RTEMS run-time implementation is based on the POSIX API. Thus the RTEMS configuration for a GNAT/RTEMS environment MUST include the --enable-

posix flag.

The following shows the command sequence required to configure, compile, and install RTEMS with the POSIX API, FreeBSD TCP/IP, and C++ support disabled. RTEMS will be built to target the BOARD_SUPPORT_PACKAGE board.

     mkdir build-rtems
     cd build-rtems
     ../rtems-4.9.1/configure --target=<TARGET_CONFIGURATION> \
         --disable-posix --disable-networking --disable-cxx \
         --enable-rtemsbsp=<BOARD_SUPPORT_PACKAGE>\
         --prefix=<INSTALL_POINT>
     make all install

Where the list of currently supported <TARGET CONFIGURATION>’s and <BOARD SUPPORT PACKAGE>’s can be found in tools/rtems-4.9.1/README.configure.

<INSTALL POINT> is typically the installation point for the tools and defaults to

/opt/rtems-4.9.1.

BSP is a supported BSP for the selected CPU family. The list of supported BSPs may be found in the file tools/rtems-4.9.1/README.configure in the RTEMS source tree. If the BSP parameter is not specified, then all supported BSPs for the selected CPU family will be built.

NOTE: The POSIX API must be enabled to use GNAT/RTEMS.

NOTE: The make utility used should be GNU make.

6 Building the Sample Applications

The RTEMS distribution includes a number of sample C, C++, Ada, and networking applications. This chapter will provide an overview of those sample applications.

6.1 Set the Environment Variable RTEMS MAKEFILE PATH

The sample application sets use the RTEMS Application Makefiles. This requires that the environment variable RTEMS_MAKEFILE_PATH point to the appropriate directory containing the installed RTEMS image built to target your particular CPU and board support package combination.

     export RTEMS_MAKEFILE_PATH=<INSTALLATION_POINT>/<CPU>-rtems/<BOARD_SUPPORT_PACKAGE>

Where <INSTALLATION POINT> and <BOARD SUPPORT PACKAGE> are those used when configuring and installing RTEMS.

NOTE: In release 4.0, BSPs were installed at <INSTALLATION_POINT>/rtems/<BOARD_ SUPPORT_PACKAGE>. This was changed to be more in compliance with GNU standards.

NOTE: GNU make is the preferred make utility. Other make implementations may work but all testing is done with GNU make.

If no errors are detected during the sample application build, it is reasonable to assume that the build of the GNU Cross Compiler Tools for RTEMS and RTEMS itself for the selected host and target combination was done properly.

6.2 Executing the Sample Applications

How each sample application executable is downloaded to your target board and executed is very dependent on the board you are using. The following is a list of commonly used BSPs classified by their RTEMS CPU family and pointers to instructions on how to use them. [NOTE: All file names should be prepended with rtems-4.9.1/c/src/lib/libbsp.] arm/edp7312 The arm/edp7312 BSP is for the ARM7-based Cogent EDP7312

board.

c4x/c4xsim The c4x/c4xsim BSP is designed to execute on any member of the Texas Instruments C3x/C4x DSP family using only on-CPU peripherals for the console and timers.

i386/pc386  See i386/pc386/HOWTO

i386/pc486 The i386/pc386 BSP specially compiled for an i486-class CPU.

i386/pc586 The i386/pc386 BSP specially compiled for a Pentium-class CPU.

i386/pc686 The i386/pc386 BSP specially compiled for a Pentium II.

i386/pck6 The i386/pc386 BSP specially compiled for an AMD K6.

m68k/gen68360 This BSP is for a MC68360 CPU. See m68k/gen68360/README for details.

m68k/mvme162  See m68k/mvme162/README.
m68k/mvme167  See m68k/mvme167/README.

mips/jmr3904 This is a BSP for the Toshiba TX3904 evaluation board simulator included with mipstx39-rtems-gdb. The BSP is located in mips/jmr3904. The TX3904 is a MIPS R3000 class CPU with serial ports and timers integrated with the processor. This BSP can be used with either real hardware or with the simulator included with mipstx39-rtems-gdb. An application can be run on the simulator by executing the following commands upon entering mipstx39-rtems-

                   gdb:
                        target sim --board=jmr3904
                        load
                        run
powerpc/mcp750  See powerpc/motorola_shared/README.
powerpc/mvme230x See powerpc/motorola_shared/README.MVME2300.

powerpc/psim This is a BSP for the PowerPC simulator included with powerpcrtems-gdb. The simulator is complicated to initialize by hand. The user is referred to the script powerpc/psim/tools/psim.

sparc/erc32 The ERC32 is a radiation hardened SPARC V7. This BSP can be used with either real ERC32 hardware or with the simulator included with sparc-rtems-gdb. An application can be run on the simulator by executing the following commands upon entering sparc-rtems-

                   gdb:
                        target sim
                        load
                        run

RTEMS has many more BSPs and new BSPs for commercial boards and CPUs with on-CPU peripherals are generally welcomed.

6.3 C/C++ Sample Applications

The C/C++ sample application set includes a number of simple applications. Some demonstrate some basic functionality in RTEMS such as writing a file, closing it, and reading it back while others can serve as starting points for RTEMS applications or libraries. Start by unarchiving them so you can peruse them. Use a command similar to the following to unarchive the sample applications:

     cd tools
     tar xjf ../archive/examples-VERSION.tgz

The sample applications most likely to be of interest to you are:

• hello world c - C Hello World application with a simple RTEMS configuration and an entry point not called main().

• simple main - Very simple program starting at main() and shutting down RTEMS via exit() without any other operations. It uses the default configuration inside RTEMS which is only intended to satisfy autoconf probes and extremely simple console-based applications.

• libcpp - Simple C++ library for RTEMS showing how to build an application library written in C++.

• psx sched report - POSIX Scheduler Reporter is a program that prints out some scheduler attributes of the RTEMS POSIX API.

Each tests is found in a separate subdirectory and built using the same command sequence. The hello_world_c sample will be used as an example.

Build the C Hello World Application

Use the following command to start the build of the sample hello world application:

     cd hello_world_c
     make

If the sample application has successfully been built, then the application executable is placed in the following directory:

     hello_world_c/o-optimize/<filename>.exe

The other C/C++ sample applications are built using a similar procedure.

6.4 Ada Sample Applications

The Ada sample application set primarily includes a a simple Hello World Ada program which can be used as a starting point for GNAT/RTEMS applications. Use the following command to unarchive the Ada sample applications:

     cd tools
     tar xjf ../archive/ada-examples-4.9.1.tar.bz2

Create a BSP Specific Makefile

Currently, the procedure for building and linking an Ada application is a bit more difficult than a C or C++ application. This is certainly an opportunity for a volunteer project.

At this time, there is a

Provided are example Makefiles for multiple BSPs. Copy one of these to the file Makefile.<BOARD SUPPORT PACKAGE> and edit it as appropriate for your local configuration.

Use the <INSTALLATION POINT> and <BOARD SUPPORT PACKAGE> specified when configuring and installing RTEMS.

6.5 Build the Sample Application

Use the following command to start the build of the sample application:

     cd tools/hello_world_ada

     make -f Makefile.<BOARD_SUPPORT_PACKAGE>

NOTE: GNU make is the preferred make utility. Other make implementations may work but all testing is done with GNU make.

If the BSP specific modifications to the Makefile were correct and no errors are detected during the sample application build, it is reasonable to assume that the build of the GNAT/RTEMS Cross Compiler Tools for RTEMS and RTEMS itself for the selected host and target combination was done properly.

6.6 Application Executable

If the sample application has successfully been build, then the application executable is placed in the following directory:

     tools/hello_world_ada/o-optimize/<filename>.exe

How this executable is downloaded to the target board is very dependent on the BOARD SUPPORT PACKAGE selected.

6.7 More Information on RTEMS Application Makefiles

These sample applications are examples of simple RTEMS applications that use the RTEMS Application Makefile system. This Makefile system simplifies building RTEMS applications by providing Makefile templates and capturing the configuration information used to build RTEMS specific to your BSP. Building an RTEMS application for different BSPs is as simple as switching the setting of RTEMS_MAKEFILE_PATH. This Makefile system is described in the

file make/README.

7 Where To Go From Here

At this point, you should have successfully installed a GNU Cross Compilation Tools for RTEMS on your host system as well as the RTEMS OS for the target host. You should have successfully linked the "hello world" program. You may even have downloaded the executable to that target and run it. What do you do next?

The answer is that it depends. You may be interested in writing an application that uses one of the multiple APIs supported by RTEMS. You may need to investigate the network or filesystem support in RTEMS. The common thread is that you are largely finished with this manual and ready to move on to others.

Whether or not you decide to dive in now and write application code or read some documentation first, this chapter is for you. The first section provides a quick roadmap of some of the RTEMS documentation. The next section provides a brief overview of the RTEMS application structure.

7.1 Documentation Overview

When writing RTEMS applications, you should find the following manuals useful because they define the calling interface to many of the services provided by RTEMS:

• RTEMS Applications C User’s Guide describes the Classic RTEMS API based on the RTEID specification.

• RTEMS POSIX API User’s Guide describes the RTEMS POSIX API that is based on the POSIX 1003.1b API.

• RTEMS ITRON 3.0 API User’s Guide describes the RTEMS implementation of the ITRON 3.0 API.

• RTEMS Network Supplement provides information on the network services provided by RTEMS.

In addition, the following manuals from the GNU Cross Compilation Toolset include information on run-time services available.

• Cygnus C Support Library describes the Standard C Library functionality provided by Newlib’s libc.

• Cygnus C Math Library describes the Standard C Math Library functionality provided by Newlib’s libm.

Finally, the RTEMS FAQ and mailing list archives are available at http://www.rtems.com.

There is a wealth of documentation available for RTEMS and the GNU tools supporting it. If you run into something that is not clear or missing, bring it to our attention.

Also, some of the RTEMS documentation is still under construction. If you would like to contribute to this effort, please contact the RTEMS Team at rtems-users@rtems.com (mailto:rtems-users@rtems.com). If you are interested in sponsoring the development of a new feature, BSP, device driver, port of an existing library, etc., please contact one of the RTEMS Service Providers listed at http://www.rtems.com/support.html

(http://www.rtems.com/support.html).

7.2 Writing an Application

From an application author’s perspective, the structure of an RTEMS application is very familiar. In POSIX language, RTEMS provides a single process, multi-threaded run-time environment. However there are two important things that are different from a standard UNIX hosted program.

First, the application developer must provide configuration information for RTEMS. This configuration information includes limits on the maximum number of various OS resources available and networking configuration among other things. See the Configuring a System in the RTEMS Applications C User’s Guide for more details.

Second, RTEMS applications may or may not start at main(). Applications begin execution at one or more user configurable application initialization tasks or threads. It is possible to configure an application to start with a single thread that whose entry point is main().

Each API supported by RTEMS (Classic, POSIX, and ITRON) allows the user to configure a set of one or more tasks that are created and started automatically during RTEMS initialization. The RTEMS Automatic Configuration Generation (confdefs.h) scheme can be used to easily generate the configuration information for an application that starts with a single initialization task. By convention, unless overridden, the default name of the initialization task varies based up API.

• Init - single Classic API Initialization Task

• POSIX_Init - single POSIX API Initialization Thread • ITRON_Init - single ITRON API Initialization Task

Regardless of the API used, when the initialization task executes, all non-networking device drivers are normally initialized, processor interrupts are enabled, and any C++ global constructors have been run. The initialization task then goes about its business of performing application specific initialization which will include initializing the networking subsystem if it is to be used. The application initialization may also involve creating tasks and other system resources such as semaphores or message queues and allocating memory. In the RTEMS examples and tests, the file init.c usually contains the initialization task. Although not required, in most of the examples, the initialization task completes by deleting itself.

As you begin to write RTEMS application code, you may be confused by the range of alternatives. Supporting multiple tasking APIs can make the choices confusing. Many application groups writing new code choose one of the APIs as their primary API and only use services from the others if nothing comparable is in their preferred one. However, the support for multiple APIs is a powerful feature when integrating code from multiple sources. You can write new code using POSIX services and still use services written in terms of the other APIs. Moreover, by adding support for yet another API, one could provide the infrastructure required to migrate from a legacy RTOS with a non-standard API to an API like POSIX.

Appendix A Using MS-Windows as a Development Host

This chapter discusses the installation of the GNU tool chain on a computer running the Microsoft Windows operating system.

This chapter was originally written by Geoffroy Montel <g montel@yahoo.com> (mailto:g_montel@yahoo.com) with input from David Fiddes <D.J@fiddes.surfaid.org> (mailto:<D.J@fiddes.surfaid.org>). It was based upon his successful but unnecessarily painful efforts with Cygwin beta versions. Cygwin and this chapter have been updated multiple times since those early days although their pioneering efforts and input is still greatly appreciated.

A.1 Microsoft Windows Version Requirements

RTEMS users report fewer problems when using Microsoft Windows NT, 2000, or XP. Although, the open source tools that are used in RTEMS development do execute on Windows 95, 98, or ME, they tend to be more stable when used with the modern Windows variants.

A.2 Cygwin

For RTEMS development, the recommended approach is to use Cygwin 1.0 or later. Cygwin is available from http://sources.redhat.com/cygwin (http://sources.redhat.com/cygwin) Recent versions of Cygwin are vastly improved over the beta versions. Most of the oddities, instabilities, and performance problems have been resolved. The installation procedure is much simpler. However, there are a handful of issues that remain to successfully use Cygwin as an RTEMS development environment.

• There is no cc program by default. The GNU configure scripts used by RTEMS require this to be present to work properly. The solution is to link gcc.exe to cc.exe as follows:

           ln -s /bin/gcc.exe  /bin/cc.exe

• Make sure /bin/sh.exe is GNU Bash. Some Cygwin versions provide a light Bourne shell which is insufficient to build RTEMS. To see which shell is installed as /bin/sh.exe, execute the command /bin/sh --version. If it looks similar to the following, then it is GNU Bash and you are OK:

           GNU bash, version 2.04.5(12)-release (i686-pc-cygwin)
           Copyright 1999 Free Software Foundation, Inc.

If you get an error or it claims to be any other shell, you need to copy it to a fake

      name and copy /bin/bash.exe to /bin/sh.exe:
           cd /bin
           mv sh.exe old_sh.exe
           cp bash.exe sh.exe

The Bourne shell has to be present in /bin directory to run shell scripts properly.

• Make sure you unarchive and build in a binary mounted filesystem (e.g. mounted with the -b option). Otherwise, many confusing errors will result.

• A user has reported that they needed to set CYGWIN=ntsec for chmod to work correctly, but had to set CYGWIN=nontsec for compile to work properly (otherwise there were complaints about permissions on a temporary file).

• If you want to build the tools from source, you have the same options as UNIX users.

• You may have to uncompress archives during this process. You must NOT use WinZip or PKZip. Instead the un-archiving process uses the GNU zip and tar programs as shown below:

     tar -xzvf archive.tgz

tar is provided with Cygwin.

A.3 Text Editor

You absolutely have to use a text editor which can save files with Unix format. So do NOT use Notepad or Wordpad! There are a number of editors freely available that can be used.

• VIM  (Vi  IMproved)  is  available  from  http://www.vim.org/

(http://www.vim.org/). This editor has the very handy ability to easily read and write files in either DOS or UNIX style.

• GNU Emacs is available for many platforms including MS-Windows. The official homepage is http://www.gnu.org/software/emacs/emacs.html

(http://www.gnu.org/software/emacs/emacs.html).  The

GNU Emacs on Windows NT and Windows 95/98 FAQ is at http://www.gnu.org/software/emacs/windows/ntemacs.html

(http://www.gnu.org/software/emacs/windows/ntemacs.html).

If you do accidentally end up with files having MS-DOS style line termination, then you may have to convert them to Unix format for some Cygwin programs to operate on them properly. The program dos2unix can be used to put them back into Unix format as shown below:

$ dos2unix XYZ
Dos2Unix: Cleaning file XYZ ...

A.4 System Requirements

Although the finished cross-compiler is fairly easy on resources, building it can take a significant amount of processing power and disk space.

• The faster the CPU, the better. The tools and Cygwin can be very CPU hungry.

• The more RAM, the better. Reports are that when building GCC and GDB, peak memory usage can exceed 256 megabytes.

• The more disk space, the better. You need more if you are building the GNU tools and the amount of disk space for binaries is obviously directly dependent upon the number of CPUs you have cross toolsets installed for.